OpenStudy (anonymous):

what is 0.68 is simple fraction form?

OpenStudy (anonymous):

i would say its 68/100

OpenStudy (anonymous):

0.68=68/100=(17*4)/(25*4)=17/25

OpenStudy (anonymous):

why is it 17x4? multiples of 68?

OpenStudy (anonymous):

yes, because you need in the simple form

OpenStudy (anonymous):

ooic..thank you

OpenStudy (anonymous):

You are welcome. It might help http://hotmath.com/hotmath_help/topics/converting-non-repeating-decimals-to-fractions.html